Let X1, X2,... ,X16 be a random sample from a N(μ, σ²) with σ² = 25 and sample mean X= 60 . Find a 95% confidence interval for μ.
1
Expert's answer
2019-04-03T11:20:27-0400

"X_1, X_2, ..., X_n" is a random sample drawn from a normal distribution with "X_i\thicksim N(\mu,\sigma^2)" for all "i". Then


"{\bar{X}-\mu \over \sigma/\sqrt{n}}\thicksim N(0,1)"

Using this probability distribution, we have 


"P(-z_{0.025}<Z<z_{0.025})=0.95"

"P(-z_{0.025}<{\bar{X}-\mu \over \sigma/\sqrt{n}}<z_{0.025})=0.95"

95% CI for µ when "\sigma^2" known and drawing from a normally distributed population:


"\bar{X}-z_{0.025}{\sigma \over \sqrt{n}}\le \mu \le \bar{X}+z_{0.025}{\sigma \over \sqrt{n}}"

"\bar{X}=60, \sigma=\sqrt {25}=5, n=16, z_{0.025}=1.96"

"60-1.96{5 \over \sqrt{16}}\le \mu \le 60+1.96{5 \over \sqrt{16}}"

"57.55\le \mu \le 62.45"